Find the area under the standard normal curve between -2.1 and 2.1.
answer:

(b) Find the area under the standard normal curve between -0.65 and 0.65.
answer:

(c) Find the area under the standard normal curve that lies to the left of 2.1.
answer:

(d) Find the area under the standard normal curve that lies to the right of 2.1.
answer:
Answer #1

a)

P( -2.1 < Z < 2.1) = P( Z < 2.1) - P( Z < -2.1)

= 0.9821 - 0.0179

= 0.9642

b)

P( -0.65 < Z < 0.65) = P( Z < 0.65) - P( Z < -0.65)

= 0.7422 - 0.2578

= 0.4844

c)

P( Z < 2.1) = 0.9821

d)

P( Z > 2.1) = 1 - P( Z < 2.1)

= 1 - 0.9821

= 0.0179